OpenStudy (anonymous):

Determine the slope and Y-intercept of the graph equation 7x + 4y = 16

OpenStudy (anonymous):

sloper intercept form is Y = mx + b B is the Y-intercept M is the slope

OpenStudy (anonymous):

if you isolate Y then you can figure them out easier 4y = -7x + 16 now divide each side by 4 and you will have the slope intercept form

OpenStudy (anonymous):

How did the 7y become a negative?

OpenStudy (anonymous):

7x + 4y = 16 we subtract 7x from BOTH sides -7x - 7x 4y = 16 - 7x

OpenStudy (anonymous):

it became a negative when she flipped the Y to other side

OpenStudy (anonymous):

@Bweezy5 ima guy

OpenStudy (anonymous):

im bad at telling guys from girls excuse me for that sir

OpenStudy (anonymous):

@Bweezy5 well, it's fine considering my pic is of two genders lol

OpenStudy (anonymous):

So Is the Y intercept 2 1/4 and the slope ??? Wait, I am really confused

OpenStudy (anonymous):

this is another graphing problem right? just want to make sure

OpenStudy (anonymous):

Yes. I am working strictly on graphing stuff right now. Chapter 3 in the math 94 class

OpenStudy (anonymous):

@ChristaB well, we would leave it as an improper fraction 7/4 but don't forget that 7 is negative

OpenStudy (anonymous):

I have Y=-7/4 + 16/4

OpenStudy (anonymous):

so the slope is -7/4 and 16/4 is the y-intercept, but simplify that xD

OpenStudy (anonymous):

16/4 simplified would be 4... So, how do I graph the equation then?

OpenStudy (anonymous):

where does it say you need to graph it? o.o

OpenStudy (anonymous):

The question reads: Determine the slope and the Y intercept of the graph of the equation. Does that not mean to graph it or does it mean to just determine the slope and Y intercept

OpenStudy (anonymous):

i dunno, didn't think u needed to graph that lol.

OpenStudy (anonymous):

i asked her and she said yea im thinking on how to graph it

OpenStudy (anonymous):

I think you are right. After I re-read the question, it doesnt say graph it

OpenStudy (anonymous):

ohh no you dont need to graph it says determine the slope not to find the slope damn i was over here pulling a muscle >.<

OpenStudy (anonymous):

 i'm guessing like that if you did need to graph anything

OpenStudy (anonymous):

Ok, that whole drawing lines on the puter thing is cool.. How do I do that?

OpenStudy (anonymous):

on the left of post there's button: Equation, Draw, and attach file

OpenStudy (anonymous):

OK... we are ready to move on to the next question. See you on the other side..lol

OpenStudy (anonymous):

im ready for you next question lol this one i was every so couldn't help sorry about that :/